OpenStudy (anonymous):

Assume a 125 pound person burns 150 calories after 15 minutes of jogging and 200 calories after 20 minutes of jogging. Part 1 (2 points): Use calories as the y-coordinate and minutes as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? Part 2 (2 points): Write an equation, in slope intercept form, to represent this data. Part 3 (2 points): How many calories will this person burn after 35 minutes of jogging? Using complete sentences, explain how the equation, slope, or graph can help to predict calories

OpenStudy (anonymous):

I just need one last help !!!

OpenStudy (anonymous):

slope m = ( 200 -150)/ ( 20 -15) = 50/5 = 10 200 = 10 * 20 + b -> b = 0 => y = 10x 35 min * 200 cal/ 20 = 350 cal